A uniform rod of length L is acted upon by two collinear forces P and Q applied to its ends in opposite directions. Calculate the tension at a distance x from the end where force P is applied.

Grade:11

1 Answers

Vikas TU
7 years ago
Net moment zero at x distance from P then,
P*x = Q*(l – x)
find x from here.
then do net torque zero at any one force and assume a T tension force at distnce x from P.
then
moment at p suppose be zero.
T*x = Q*l
T = Ql/x
and put x value from the above calculated.